2x-y = -4 -3x+y=-9
Answer by saifoo.khan(162)

2x-y = -4 -3x+y=-9
Hello,
x=13
y=30
If u want the steps, i can provide u as well.
THanks.
Saifoo.khan
einstein.saifoo@gmail.com